# Finitely supertranslated Schwarzschild black hole and its perturbations

### v2

@@ I Introduction
The search for memory effects has been conducted, but no strong evidence has been found by LIGO-Virgo-KAGRA collaboration [35, 36, 37, 38, 39, 40, 41] or the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [42]. The prospects of its future detection have been anticipated by several works. It was found out that with the networks of the ground-based interferometers, one may observe the memory effect by coherently adding the signals from hundreds or thousands stellar mass binary black hole coalescences after 3 - 5 years operation [43, 37, 44]. TianQin [45, 46] and LISA [47, 48] are able to detect the memory signals produced by massive binary black hole mergers.
@@
In fact, the physical influence of a supertranslated black hole was already considered previously. For example, Hawking, Perry and Strominger [49] argued the soft hair of a supertranslated Vaidya black hole [52] may resolve the information paradox [53]. The Hawking radiation of this black hole was later derived with the tunneling method in Ref. [54], and its spectrum depends on the soft hair. The Hawking radiation of a supertranslated Schwarzschild black hole was also studied in Ref. [55], and the spectrum is not altered. In these studies, an infinitesimally supertranslated black hole is considered. In the current work, the focus will be on a finitely supertranslated Schwarzschild black hole. One shall note that an infinitesimally supertranslated metric differs from the original one by a small perturbation, while a finitely supertranslated one differs by an arbitrarily large amount. Since the gravitation is nonlinear, it is interesting to study the perturbations of a finitely supertranslated black hole.
@@
Instead, one shall take advantage of the general covariance of the Einstein’s equation. That is, if a solution to Einstein’s equation is determined in one coordinate system, the transformed solution under a diffeomorphism solves the transformed Einstein’s equation. Therefore, the strategy here is to take the standard perturbative solutions of the Schwarzschild spacetime and transform them to a suitable Bondi-Sachs coordinates, in which the background metric describes the supertranslated black hole. Then, one obtains the supertranslated perturbations. One cannot simply apply the coordinate transformation provided by Ref. [56], as it generally does not transform the perturbed Schwarzschild metric to the desired form. And even worse, one cannot easily generalize the method in Ref. [56] to get a new coordinate transformation for the perturbed metric. The very coordinate transformation derived in that work heavily relies on the fact that the background metric is time-independent, while the perturbed metric is usually not.
Since we are eventually interested in detecting the presence of a supertranslated black hole by observing its perturbations, we merely have to perform a suitable coordinate transformation near the null infinity, where the measurement is conducted. Such a finite supertranslation has already been derived in Ref. [57] for a generic asymptotically flat spacetime. Applying this finite supertranslation transform would give the perturbations for the supertranslated Schwarzschild black hole. It is thus found out that the supertranslated perturbation can be easily obtained by shifting the retarded time $u$ of the standard results by a function of angles, $\beta(\theta,\phi)$, which quantifies the supertranslation. Like Eq. (1), the supertranslated metric obtained with a finite supertranslation is not invariant under the parity transformation. In fact, it might not be sensible to have some transformation named parity. Nevertheless, by studying the property of the parity transformation in the usual Schwarzschild spacetime, we proposed a new transformation named asymptotic parity transformation. Under this transformation, it is meaningful to state that the spectra of the odd and even parity perturbations are the same as the standard ones, and the isospectrality is not violated.
@@
This work is organized in the following way. After the presentation of the notations and conventions used in this work in Sec. I.1, some basics of the asymptotically flat spacetime will be reviewed in Sec. II. The finite supertranslation will be discussed in Sec. II.1. Then, Sec. III will be devoted to the discussion of the standard Schwarzschild metric (Sec. III.1) and the supertranslated one (Sec. III.2). Section IV reviewed the black hole perturbation theory, in particular, the gauge-invariant formalism. After that, the standard black hole perturbations are expressed in the suitable Bondi-Sachs coordinates in Sec. V. Since there are two types of perturbations, we start with the even parity perturbation in Sec. V.1, and then the odd parity perturbation in Sec. V.2. Section V.3 is a short intermediate summary. The supertranslated perturbations are obtained in Sec. VI, and their spherical decomposition is discussed in Sec. VI.1. In Section VII, we speculate the possible ways to detect the supertranslated perturbation. Finally, there is a conclusion VIII. In Appendix A, the generalized angular momentum operators are defined for tensors on the unit 2-sphere.
@@ I.1 Notations and conventions
In the unhatted coordinates, the metric is denoted by $g_{\mu\nu}$ and its line element by $\mathrm{d}s^{2}$, while in the hatted coordinates, they are $\hat{g}_{\mu\nu}$ and $\mathrm{d}\hat{s}^{2}$. For the background metric, its element is $\mathrm{d}\mathring{s}^{2}$ in the unhatted coordinates and $\mathrm{d}\hat{\mathring{s}}^{2}$ in the hatted ones. Like the line element, many quantities will be hatted if they are expressed in the hatted coordinates. The round metric on a unit 2-sphere is $\gamma_{AB}$, which is $\gamma_{AB}\mathrm{d}\theta^{A}\mathrm{d}\theta^{B}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}$. Its inverse is $\gamma^{AB}$, volume element $\epsilon_{AB}$, determinant $\gamma$, and covariant derivative $\mathscr{D}_{A}$. The capital Latin indices will be raised or lowered with $\gamma^{AB}$ or $\gamma_{AB}$, respectively. For the standard Schwarzschild metric, its temporal-radial components are labeled by $\bar{g}_{ab}$, whose inverse is $\bar{g}^{ab}$, volume element $\bar{\epsilon}_{ab}$ and covariant derivative ¯$\bar{\nabla}_{a}$. The lower case Latin indices will be naturally raised or lowered with $\bar{g}^{ab}$ or $\bar{g}_{ab}$, respectively.
@@ II Asymptotically flat spacetimes
with $Y^{\ell m}$ the spherical harmonics, and the last part represents a proper supertranslation. Note that the first two terms are actually the linear combination of $Y^{\ell m}$ with $\ell=0,1$. $\mathcal{Y}^{A}$ can be written as
@@
with $\omega^{\mu\nu}$ an antisymmetric constant tensor. The first term generates the Lorentz boosts, and the second the spatial rotation. As discussed in Introduction, super-Lorentz transformations are allowed in the extended or generalized BMS Groups. Their generators are given by more general expressions. For example, in the generalized BMS group, $\mathcal{Y}^{A}=\mathscr{D}^{A}Y^{\ell m}+\epsilon^{AB}\mathscr{D}_{B}Y^{\ell m}$.
@@ III.1 The standard Schwarzschild black hole
The coordinate systems in Eq. (19) are adopted to the symmetries of the metric. This results in important consequences. In particular, for perturbations about the standard Schwarzschild metric, their equations can be simplified a lot, and the solutions can be derived with the method of the separation of variables. For example, for a massless scalar perturbation $S$, the Klein-Gordon equation is
@@ VI Supertranslated perturbations
The supertranslated bases have the following interesting properties. First, $c_{AB}^{(e,n\ell)}(u+\beta)$ and $c_{AB}^{(o,n\ell)}(u+\beta)$ are eigenfunctions of $\hat{\chi}_{t}$ with eigenvalues $i\omega_{n\ell}$, which can be verified directly. Second, they are also eigenfunctions of the operators $\hat{\mathscr{D}}^{2}$ and $\hat{\chi}_{z}$ with eigenvalues $4-\ell(\ell+1)$ and $im=0$, respectively. Again, please refer to Appendix A for the definitions of these operators. Finally, they have well-defined parities under $\hat{\mathscr{P}}$, defined in Eq. (30). Indeed, under $\hat{\mathscr{P}}$, one finds out that
@@ VII Speculation on detection
are invariant. These differences depend on the background spacetime. If one performs another proper supertranslation $\beta^{\prime}(\bm{\theta})$, the background supertranslated metric changes, with the new shear tensor $\hat{\mathring{c}}^{\prime}_{AB}=\hat{\mathring{c}}_{AB}-2\mathscr{D}_{\langle A}\mathscr{D}_{B\rangle}\beta^{\prime}$. At the same time, the phase shifts become $\delta\phi^{\prime(k)}_{02}=\delta\phi^{(k)}_{02}+\varpi_{02}\beta^{\prime}(\bm{\theta}_{k})$. So
@@
The second line does not vanish in general. These differences truly characterizes the supertranslated spacetime.
@@ VIII Conclusion
We started with the supertranslated Schwarzschild black hole, due to its simplicity. Although Ref. [56] already derived the closed form metric for such a black hole, we did not work out its perturbation equation due to its extreme complexcity and its lack of the apparent spherical symmetry. Instead, we made use of the general covariance of Einstein’s equation, and obtained the quasinormal modes by suitably supertranslating the known perturbations of the standard Schwarzschild black hole. This method can be easily generalized to the Kerr black hole.
@@
With the gravitational wave interferometers, one may measure the phase shift proportional to $\beta(\bm{\theta})$ of the gravitational wave strain during the ringdown stage. Due to the freedom of performing the time translation, one shall use multiple detectors, and if the differences in the phase shifts are nonvanishing, one may infer the possible existence of the supertranslated black hole. A single detector may also achieve this goal if the supertranslated quasinormal mode is gravitational lensed.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grant Nos. 11633001 and 11920101003, the Key Program of the National Natural Science Foundation of China under grant No. 12433001, and the Strategic Priority Research Program of the Chinese Academy of Sciences, grant No. XDB23000000. S. H. was supported by the National Natural Science Foundation of China under Grant No. 12205222. K.L. was supported by Fapesq-PB of Brazil.
@@ Appendix A Generalizing the angular momentum operators
In the treatment of the black hole perturbations of the standard Schwarzschild black hole, the perturbations are conveniently written as the linear combinations of eigenfunctions of operators $\partial_{u}$, $\mathscr{D}^{2}$, $\chi_{i}$’s and $\mathscr{P}$ due to the nice symmetries possessed by the background metric. For the supertranslated black hole, the metric is not explicitly invariant under the spatial rotations. As discussed previously, the supertranslated metric is still implicitly invariant under the transformations generated by $\hat{\chi}_{i}$’s. Although the perturbation to the supertranslated black hole may not be decomposed into the eigenfunctions of $\mathscr{D}^{2}$ and $\chi_{i}$’s, it is still possible to find the generalizations of these operators such that the perturbation is the linear combination of their eigenfunctions.
@@
and a similar equation for $X_{AB}^{\ell m}$ can also be obtained. Here, one shall repeatedly use Eqs. (91a) and (89). Comparing the above result with Table 1, one knows that the defintion (90) works.

### v3

@@ I Introduction
The search for memory effects has been conducted, but no strong evidence has been found by LIGO-Virgo-KAGRA collaboration [35, 36, 37, 38, 39, 40, 41] or the North American Nanohertz Observatory for Gravitational Waves (NANOGrav) [42]. The prospects of its future detection have been anticipated by several works. It was found out that with the networks of the ground-based interferometers, one may observe the memory effect by coherently adding the signals from hundreds or thousands stellar mass binary black hole merger events after 3 - 5 years operation [43, 37, 44]. TianQin [45, 46] and LISA [47, 48] are able to detect the memory signals produced by massive binary black hole mergers. Notably, it was recently predicted that DECIGO could detect a few thousand strong enough memory signals produced by stellar-mass binary black holes [49].
@@
In fact, the physical influence of a supertranslated black hole was already considered previously. For example, Hawking, Perry and Strominger [50] argued the soft hair of a supertranslated Vaidya black hole [53] may resolve the information paradox [54]. The Hawking radiation of this black hole was later derived with the tunneling method in Ref. [55], and its spectrum depends on the soft hair. The Hawking radiation of a supertranslated Schwarzschild black hole was also studied in Ref. [56], and the spectrum is not altered. In these studies, the infinitesimally supertranslated black holes were considered. In the current work, the focus will be on a finitely supertranslated Schwarzschild black hole. One shall note that an infinitesimally supertranslated metric differs from the original one by a small perturbation, while a finitely supertranslated one differs by an arbitrarily large amount. Since the gravitation is nonlinear, it is interesting to study the perturbations of a finitely supertranslated black hole.
@@
Instead, one shall take advantage of the general covariance of the Einstein’s equation. That is, if a solution to Einstein’s equation is determined in one coordinate system, the transformed solution under a diffeomorphism solves the transformed Einstein’s equation. Therefore, the strategy here is to take the standard perturbative solutions of the Schwarzschild spacetime and transform them to the suitable Bondi-Sachs coordinates, in which the background metric describes the supertranslated black hole. Then, one obtains the supertranslated perturbations. One cannot simply apply the coordinate transformation provided by Ref. [57], as it generally does not transform the perturbed Schwarzschild metric to the desired form. And even worse, one cannot easily generalize the method in Ref. [57] to get a new coordinate transformation for the perturbed metric. The very coordinate transformation derived in that work heavily relies on the fact that the background metric is time-independent, while the perturbed metric is usually not.
Since we will be eventually interested in detecting the presence of a supertranslated black hole by observing its perturbations, we merely have to perform a suitable coordinate transformation near the null infinity, where the measurement is conducted. Such a finite supertranslation has already been derived in Ref. [58] for a generic asymptotically flat spacetime. Applying this finite supertranslation transformation would give the perturbations for the supertranslated Schwarzschild black hole. It was thus found out that the supertranslated perturbation can be easily obtained by shifting the retarded time $u$ of the standard results by a function of angles, $\beta(\theta,\phi)$, which quantifies the supertranslation. Like Eq. (1), the supertranslated metric obtained with a finite supertranslation is not invariant under the parity transformation. In fact, it might not be sensible to have some transformation named parity. Nevertheless, by studying the property of the parity transformation in the usual Schwarzschild spacetime, we proposed a new transformation named asymptotic parity transformation. Under this transformation, it is meaningful to state that the spectra of the odd and even parity perturbations are the same as the standard ones, and the isospectrality is not violated.
@@
This work is organized in the following way. After the presentation of the notations and conventions used in this work in Sec. I.1, some basics of the asymptotically flat spacetime will be reviewed in Sec. II. The finite supertranslation will be discussed in Sec. II.1. Then, Sec. III will be devoted to the discussion of the standard Schwarzschild metric (Sec. III.1) and the supertranslated one (Sec. III.2). Section IV reviewed the black hole perturbation theory, in particular, the gauge-invariant formalism. After that, the standard black hole perturbations are expressed in the suitable Bondi-Sachs coordinates in Sec. V. Since there are two types of perturbations, we start with the even parity perturbation in Sec. V.1, and then the odd parity perturbation in Sec. V.2. Section V.3 is a short intermediate summary. The supertranslated perturbations are obtained in Sec. VI, and their spherical decomposition is discussed in Sec. VI.1. In Section VII, we speculate the possible ways to detect the supertranslated perturbation. Finally, there is a conclusion in VIII. In Appendix A, the generalized angular momentum operators are defined for tensors on the unit 2-sphere. In Appendix B, the covariant boundary conditions for the quasinormal modes are discussed.
@@ I.1 Notations and conventions
In the unhatted coordinates, the metric is denoted by $g_{\mu\nu}$ and its line element by $\mathrm{d}s^{2}$, while in the hatted coordinates, they are $\hat{g}_{\mu\nu}$ and $\mathrm{d}\hat{s}^{2}$, respectively. For the background metric, its line element is $\mathrm{d}\mathring{s}^{2}$ in the unhatted coordinates and $\mathrm{d}\hat{\mathring{s}}^{2}$ in the hatted ones. Like the line element, many quantities will be hatted if they are expressed in the hatted coordinates. The round metric on a unit 2-sphere is $\gamma_{AB}$, which is $\gamma_{AB}\mathrm{d}\theta^{A}\mathrm{d}\theta^{B}=\mathrm{d}\theta^{2}+\sin^{2}\theta\mathrm{d}\phi^{2}$. Its inverse is $\gamma^{AB}$, volume element $\epsilon_{AB}$, determinant $\gamma$, and covariant derivative $\mathscr{D}_{A}$. The capital Latin indices will be raised or lowered with $\gamma^{AB}$ or $\gamma_{AB}$, respectively. For the standard Schwarzschild metric, its temporal-radial components are labeled by $\bar{g}_{ab}$, whose inverse is $\bar{g}^{ab}$, volume element $\bar{\epsilon}_{ab}$ and covariant derivative ¯$\bar{\nabla}_{a}$. The lower case Latin indices will be naturally raised or lowered with $\bar{g}^{ab}$ or $\bar{g}_{ab}$, respectively.
@@ II Asymptotically flat spacetimes
with $Y^{\ell m}$ the spherical harmonics, and the last part represents a proper supertranslation. Note that the first two terms are actually the linear combinations of $Y^{\ell m}$ with $\ell=0,1$. $\mathcal{Y}^{A}$ can be written as
@@
with $\omega^{\mu\nu}$ an antisymmetric constant tensor. The first term generates the Lorentz boost, and the second the spatial rotation. As discussed in the Introduction, super-Lorentz transformations are allowed in the extended or generalized BMS groups. Their generators are given by more general expressions. For example, in the generalized BMS group, $\mathcal{Y}^{A}=\mathscr{D}^{A}Y^{\ell m}+\epsilon^{AB}\mathscr{D}_{B}Y^{\ell m}$.
@@ III.1 The standard Schwarzschild black hole
The coordinate systems in Eq. (19) are adopted to the symmetries of the metric. This results in some important consequences. In particular, for perturbations about the standard Schwarzschild metric, their equations can be simplified a lot, and the solutions can be derived with the method of the separation of variables. For example, for a massless scalar perturbation $S$, the Klein-Gordon equation is
@@ VI Supertranslated perturbations
The supertranslated bases have the following interesting properties. First, $c_{AB}^{(e,n\ell)}(u+\beta)$ and $c_{AB}^{(o,n\ell)}(u+\beta)$ are eigenfunctions of $\hat{\chi}_{t}$ with eigenvalues $i\omega_{n\ell}$, which can be verified directly. Second, they are also eigenfunctions of the operators $\hat{\mathscr{D}}^{2}$ and $\hat{\chi}_{z}$ with eigenvalues $4-\ell(\ell+1)$ and $im=0$, respectively. Again, please refer to Appendix A for the definitions of these operators. Third, they satisfy the correct boundary condition at the infinity. Here, we only present the supertranslated perturbation near the infinity, so one shall check whether the complete perturbation satisfies the boundary condition near the horizon. This is done in Appendix B, where certain covariant boundary conditions, equivalent to Eq. (50), are proposed. Finally, they have well-defined parities under $\hat{\mathscr{P}}$, defined in Eq. (30). Indeed, under $\hat{\mathscr{P}}$, one finds out that
@@ VII Speculation on detection
are invariant. We emphasize that one shall shift all time axes for these interferometers by the same amount, otherwise, the time shift would be angular dependent, $u\rightarrow u+\beta^{\prime}(\bm{\theta})$. $\beta^{\prime}(\bm{\theta})$ may actually parameterize another supertranslation, and this changes the background spacetime metric, which is given by Eq. (26) with $\beta$ replaced by $\beta+\beta^{\prime}$. In particular, the shear tensor changes, $\hat{\mathring{c}}^{\prime}_{AB}=\hat{\mathring{c}}_{AB}-2\mathscr{D}_{\langle A}\mathscr{D}_{B\rangle}\beta^{\prime}$. This is not the metric that one started with. Since we are interested in the quasinormal modes of the original metric, it is not permissible to shift the time axes of the interferometers independently. Of course, if one insists on shifting the time axes independently, one actually measures the quasinormal modes for a different supertranslated Schwarzschild black hole, and the phase shifts become $\delta\phi^{\prime(k)}_{02}=\delta\phi^{(k)}_{02}+\varpi_{02}\beta^{\prime}(\bm{\theta}_{k})$. So
@@
The second line does not vanish in general. These phase differences depend on the supertranslation, and truly characterize the supertranslated spacetime.
@@ VIII Conclusion
We started with the supertranslated Schwarzschild black hole, due to its simplicity. Although Ref. [57] already derived the closed form metric for such a black hole, we did not work out its perturbation equation due to its extreme complexity and its lack of the apparent spherical symmetry. Instead, we made use of the general covariance of Einstein’s equation, and obtained the quasinormal modes by suitably supertranslating the known perturbations of the standard Schwarzschild black hole. This method can be easily generalized to the Kerr black hole.
@@
With the gravitational wave interferometers, one may measure the phase shift proportional to $\beta(\bm{\theta})$ of the gravitational wave strain during the ringdown stage. Due to the freedom of performing the time translation, one shall use multiple detectors, and if the differences in the phase shifts are nonvanishing, one may infer the possible existence of the supertranslated black hole. A single detector may also achieve this goal if the supertranslated quasinormal mode is gravitationally lensed.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No. 11633001 and No. 11920101003, the Key Program of the National Natural Science Foundation of China under Grant No. 12433001, and the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB23000000. S. H. was supported by the National Natural Science Foundation of China under Grant No. 12205222. K.L. was supported by Fapesq-PB of Brazil.
@@ Appendix A Generalizing the angular momentum operators
In the treatment of the black hole perturbations of the standard Schwarzschild black hole, the perturbations are conveniently written as the linear combinations of eigenfunctions of operators $\partial_{u}$, $\mathscr{D}^{2}$, $\chi_{i}$’s, and $\mathscr{P}$ due to the nice symmetries possessed by the background metric. For the supertranslated black hole, the metric is not explicitly invariant under the spatial rotations. As discussed previously, the supertranslated metric is still implicitly invariant under the transformations generated by $\hat{\chi}_{i}$’s. Although the perturbation to the supertranslated black hole may not be decomposed into the eigenfunctions of $\mathscr{D}^{2}$ and $\chi_{i}$’s, it is still possible to find the generalizations of these operators such that the perturbation is the linear combination of their eigenfunctions.
@@
and a similar equation for $X_{AB}^{\ell m}$ can also be obtained. Here, one shall repeatedly use Eqs. (91a) and (89). Comparing the above result with Table 1, one knows that the definition (90) works.
@@
Appendix B The covariant boundary conditions
Equation (74) shows the supertranslated bases for quasinormal modes. One concludes that both $c^{(e,n\ell)}_{AB}(u+\beta)$ and $c^{(o,n\ell)}_{AB}(u+\beta)$ are outgoing, satisfying the boundary condition of the quasinormal mode at the infinity. In this appendix, we will show that the boundary conditions at the horizon and the infinity are both satisfied after an arbitrary coordinate transformation.
For that purpose, one shall reexpress the boundary condition in a covariant manner. The boundary condition given by (50) actually states that at both the horizon and the infinity, the gravitational energy flux shall propagate into the horizon and the infinity. Given the phase of the gravitational wave $\Phi$, the direction of the flux can be represented by
$k^{\mu}=-\nabla^{\mu}\Phi,$ (100)
near the horizon and the infinity, according to the short-wavelength approximation [65, 108]. Note that near the horizon, the wave is highly blueshifted relative to the local inertia observer, and near the infinity, the background spacetime is basically Minkowski. So the short-wavelength approximation is satisfied.
Next, let us define a level function $\Omega$, taking the advantage of the spherical symmetry of the Schwarzschild spacetime. With this level function $\Omega$, the spacetime is foliated into leaves, which are sets of the orbits of the isometry group SO(3) [12]. The horizon occurs at $\Omega_{H}$, and as one moves away from the horizon to the infinity, $\Omega$ increases monotonically and becomes infinite at the infinity. Note that the Schwarzschild spacetime is always spherically symmetry even if one does not use the (standard) Schwarzschild coordinates to describe it. In a coordinate system adapted to the symmetry, the metric is explicitly spherically symmetric, while in a coordinate system not compatible to the symmetry, the metric is implicitly spherically symmetric. So in any coordinate system, one can always define such a level function $\Omega$ and foliate the spacetime accordingly.
Now, with the wave vector $k^{\mu}$ and the level function $\Omega$, one can write down the boundary condition at the horizon
$\lim_{\Omega\rightarrow\Omega_{H}}k^{\mu}\nabla_{\mu}\Omega<0.$ (101)
Near the infinity, one has
$\lim_{\Omega\rightarrow\infty}k^{\mu}\nabla_{\mu}\Omega>0,$ (102)
instead.
One can show that Eq. (50) agrees with these covariant boundary conditions. In Eq. (50), the adapted coordinate system is used, so the level function can be taken to be $\Omega=r$. Near the infinity, the phase of the wave is $\Phi=\varpi u\approx\varpi(t-r)$ with $\varpi$ the real frequency, so the wave vector is $k^{\mu}\approx-\varpi(\delta^{\mu}_{t}-\delta^{\mu}_{r})$, and $\lim_{\Omega\rightarrow\infty}k^{\mu}\nabla_{\mu}\phi=\varpi>0$. Near the horizon, the phase is given by $\Phi=\varpi v=\varpi(t+r_{*})$, so one has $k^{\mu}=-\varpi(g^{\mu t}+g^{\mu r}\mathrm{d}r_{*}/\mathrm{d}r)$, and $\lim_{\Omega\rightarrow\Omega_{H}}k^{\mu}\nabla\Omega=\lim_{\Omega\rightarrow\Omega_{H}}-\varpi g^{rr}\mathrm{d}r_{*}/\mathrm{d}r=-\varpi<0$.
On the other hand, one can use the covariant boundary conditions to reproduce Eq. (50). Indeed, for any phase $\Phi$, one has $\nabla_{\mu}\Phi=\Phi_{t}\nabla_{\mu}t+\Phi_{r}\nabla_{\mu}r$. It has no components in the angular directions because the angular dependence is encoded in the tensor spherical harmonics. Near the infinity, $\lim_{\Omega\rightarrow\infty}k^{\mu}\nabla_{\mu}\Omega=-\Phi_{r}|_{\infty}>0$, and the null condition $k^{\mu}k_{\mu}=0$ implies that $\Phi_{t}|_{\infty}=-\Phi_{r}|_{\infty}$. The negative sign is due to the fact that $k^{\mu}$ shall be future-pointing. So $\nabla_{\mu}\Phi|_{\infty}=\Phi_{t}|_{\infty}\nabla_{\mu}(t-r)=\Phi_{t}|_{\infty}\nabla_{\mu}u$, and $\Phi|_{\infty}=\Phi_{t}|_{\infty}u$, modulo a constant phase. Near the horizon, $\lim_{\Omega\rightarrow\Omega_{H}}k^{\mu}\nabla_{\mu}\Omega=-\lim_{r\rightarrow r_{H}}\Phi_{r}g^{rr}<0$, and $\Phi_{t}=\Phi_{r}g^{rr}$ due to the null condition. So, $\nabla_{\mu}\Phi=\Phi_{r}g^{rr}(\nabla_{\mu}t+g_{rr}\nabla_{\mu}r)=\Phi_{r}g^{rr}(\nabla_{\mu}t+\nabla_{\mu}r_{*})=\Phi_{r}g^{rr}\nabla_{\mu}v$, which implies that $\lim_{\Omega\rightarrow\Omega_{H}}\Phi=\lim_{r\rightarrow r_{H}}(\Phi_{r}g^{rr})v$.
In the end, one may generalize the above covariant boundary conditions to more general level functions $\Omega^{\prime}$, as long as
$\Omega^{\prime}=\left\{\begin{array}[]{cc}\Omega+(\Omega-\Omega_{H})a+O(\Omega-\Omega_{H})^{2}&,\Omega\rightarrow\Omega_{H},\\ \Omega+O(1/\Omega)&,\Omega\rightarrow\infty,\end{array}\right.$
where $a>-1$ is independent of $\Omega$.